OpenStudy (anonymous):

it takes Brenda 7 hours to oil the lanes in a bowling alley. gabriella can oil the same lanes in 6 hours. how long would it take them if they worked together?

OpenStudy (amistre64):

these problems are always easiest when you can figure out how much work is done by each person in "1" hour...

OpenStudy (amistre64):

If B takes 7 hours to complete it by herself; then in 1 hour she does 1/7 of the work if G takes 6 hours to complete it by herself; then in 1 hour she does 1/6 of the work. Working together; they should do: 1/6 + 1/7 of the work in 1 hour.

OpenStudy (amistre64):

13/42 get done in one hour if they work together. 13/42 times how many hours equals 42/42?

OpenStudy (anonymous):

The actual formula is (x/the amount of time for one person) + (x/the amount of time for the second person) = 1 so we would have x/7 + x/6 = 1 Common denominator of 42 should be multiplied by each term 42(x/7) + 42(x/6) = 42(1) cancelling you would get 6x + 7x = 42 13x = 42 divide by 13 x = 4 It would take them 4 hours working together.

OpenStudy (amistre64):

1 hour = 13/42 2 hours = 26/42 3 hours = 39/42 3 hours and something so far :)

OpenStudy (anonymous):

o okay thank u

OpenStudy (amistre64):

4(13) = 52, not 42....

OpenStudy (amistre64):

3 hours and 1/14th of another.... which is less than 5 minutes